# Optical-field-induced dips and splits in nonlinear spectra of selective reflection from high-density atomic vapor

### v1

@@ 3 Experimental results and discussions
$N_{i}$, $10^{17}$ cm-3 Slope $a$ Intercept $b$, GHz
3.6 $0.93\pm 0.16$ $0.20\pm 0.95$
2.5 $0.59\pm 0.08$ $1.73\pm 0.48$
1.7 $0.42\pm 0.03$ $2.74\pm 0.20$
1.2 $\pm 0.02 3.01±0.13$$\pm 0.1$
The slopes of fitted lines vs density are presented in the inset in Fig. 4. The theoretical slope, obtained by the expression (2), is equal to ”unity”. The difference between the ”unity” and the slope of the fitted line for the maximal density $N_{4}$ is near 7%. This small difference supports our interpretation of the experimental results by ”dressed atomic states” description CohenTannoudjibook1992 , Boydbook2007 . In Table 1 the small values of the slope a at the densities $N_{1}$ and $N_{2}$ reveal the inhomogeneous character of the line-shapes. The frequency intervals $\Delta\nu_{\text{SP}}$ in Fig. 4 are measured between maximums of spectral dependences of derivative $\text{d}R/\text{d}\nu$, separated by the optically saturated dips. The corresponding intercepts $b$ at $N_{1}$ and $N_{2}$ in Table 1 can be associated with homogeneous part of the widths. The intermediate values of $a$ and $b$, measured at the density $N_{3}$, may be explained by a transition from the inhomogeneous self-broadening to the homogeneous self-broadening in this density region. The transition may be a reason for a possible reduction of the contribution of the hfs-components to spectral shapes of the symmetrical resonances in Fig. 3. May be there is an analogy between the reduction of hfs-components in rubidium resonance transition due to the density variation (Fig. 3) and the density dependence of Zeeman splitting of D-lines in potassium vapor VASZeemanPRA1997 . Probably these effects may be associated with many-body interactions in high-density atomic gases.
@@ 4 Conclusions
This paper presents the results of an experimental study of nonlinear selective reflection from high-density rubidium atomic vapor, where self-broadening of the resonance transition dominates over Doppler width. In the experiment the pump-probe spectroscopic technique with two lasers are used. The reflection spectra of weak probe beam are recorded at several atomic densities in the range $(1.2\text{--}3.6)\cdot 10^{17}$ cm-3 and various pump beam intensities. With growing density, the broadening character was changed from inhomogeneous to homogeneous. The strong pump field splits the spectra on two resonances. At maximal density, these homogeneously broadened symmetrical resonances are results of the Rabi oscillations Demtroderbook2008 . The appearance of the symmetrical resonances can be explained more strictly in the frame of ”dressed atomic states” approach CohenTannoudjibook1992 , Boydbook2007 . At the maximal density the experimental conditions are appropriated for research of the dressed atomic states. Now we can say that the main goal of our research is reached. At lower densities the spectral profiles are inhomogeneously broadened. Spectral profiles of the frequency derivative are split by the optically saturated dips. At low densities the broadening of spectral profiles is inhomogeneous. Spectral profiles of the frequency derivative are split by the optically saturated dips. The width of such dips is a combination of the homogeneous part of self-broadening and intensity dependent field broadening. Additional studies of the transition from inhomogeneous to homogeneous broadening may initiate a further development of a theory of interatomic interactions in a high-density atomic gas.
In the nanocells SarkisyanPRL2012 , SarkisyanPRL2018 the dressed atomic states can be investigated in absorption spectra of the dense atomic gases. The unique nonlinear properties of ultrathin layers of ”dressed dense atomic gases” may be useful for the development of new quantum devices. The ”dressed states” spectroscopy may be applied for a study of many-body interactions in ultracold nonideal plasmas with the resonance transitions in the visible range ZelenerJETPL2021 , ZelenerPRL2024 .
@@
Vladimir Sautenkov: Conceptualization, Methodology, Validation, Writing – original draft, Funding acquisition. Sergey Saakyan: Investigation, Software, Validation, Formal analysis, Writing – review & editing. Andrei Bobrov: Formal analysis, Methodology, Writing – review & editing. Boris B. Zelener: Supervision, Resources, Writing – review & editing, Funding acquisition.

### v2

@@ 3 Experimental results and discussions
$i$ $N_{i}$, $10^{17}$ cm-3 Slope $a_{i}$ Intercept $b_{i}$, GHz
1 $1.2$ $\pm 0.02 3.01±0.13$$\pm 0.1$
2 $1.7$ $0.42\pm 0.03$ $2.74\pm 0.20$
3 $2.5$ $0.59\pm 0.08$ $1.73\pm 0.48$
4 $3.6$ $0.93\pm 0.16$ $0.20\pm 0.95$
The slopes of the fitted lines versus density are presented in the inset in Fig. 4. The theoretical slope, obtained by the expression (2), is equal to unity. The difference between the ”unity” and the slope of the fitted line for the highest density $N_{4}$ is near 7%. This small difference supports our interpretation of the experimental results as ”dressed atomic states” CohenTannoudjibook1992, Boydbook2007. In Table 1 the small values of the slope $a$ at the densities $N_{1}$ and $N_{2}$ reveal the inhomogeneous character of the line shapes. The frequency intervals $\Delta\nu_{\text{SP}}$ in Fig. 4 are measured between the maxima of the spectral dependences of the derivative $\text{d}R/\text{d}\nu$, separated by the optically saturated dips. The corresponding intercepts $b_{i}$ at $N_{1}$ and $N_{2}$ in Table 1 can be associated with the homogeneous part of the widths. The intermediate values of $a_{3}$ and $b_{3}$, measured at the density $N_{3}$, can be explained by a transition from the inhomogeneous self-broadening to the homogeneous self-broadening in this density region. The transition from the inhomogeneous self-broadening (densities $N_{1}$, $N_{2}$ and $N_{3}$) to the homogeneous (density $N_{4}$) can be attributed to the quasi-molecular effects in this density region VASZeemanPRA1997. The quasi-molecular regime can appear in a density range where inter-atomic interactions are binary but the symmetry of the atomic wave function is changed due to the rapid succession of two-particle collisions. As a result, the atomic angular momenta could be modified and hfs components in atomic spectra could be reduced. We see an analogy between the observed reduction of hfs components in rubidium spectra (Fig. 3) and the dependence of Zeeman splitting of $D$-lines in potassium vapor on the atomic density VASZeemanPRA1997. It shall be noted that in our experiment a contribution of the quadrupole-quadrupole interaction from 5P5P5P-5P coupling to selective reflection spectra were not observed. Our data are consistent with the theoretical results of Ref. MukamelPRA1994, where higher multipole terms were neglected (e.g. the quadrupole-quadrupole interaction from the 5P5P5P-5P coupling). The calculations in Ref. MukamelPRA1994 were performed for atomic density range in which an average distance between atoms is much larger than the Weisskopf radius.
@@ 4 Conclusions
This paper presents the results of an experimental study of nonlinear selective reflection from high-density rubidium atomic vapor, where self-broadening of the resonance transition dominates over the Doppler width. In the experiment the pump-probe spectroscopic technique with two lasers are used. The reflection spectra of weak probe beam are recorded at several atomic densities in the range $(1.2\text{--}3.6)\times 10^{17}$ cm-3 and various pump beam intensities. As the density increased, the broadening mechanism transitioned from inhomogeneous to homogeneous. The strong pump field split the spectra into two resonances. At the highest density these homogeneously broadened symmetric resonances are result from the Rabi flopping oscillations Demtroderbook2008. The appearance of the symmetric resonances can be explained more strictly within the framework of ”dressed atomic states” approach CohenTannoudjibook1992, Boydbook2007. At the highest density the experimental conditions are appropriated for research of the dressed atomic states. Now we can say that the main goal of our research is reached. At lower densities the spectral profiles are inhomogeneously broadened. Spectral profiles of the frequency derivative are separated by optically saturated dips. The width of such dips is a combination of the homogeneous part of self-broadening and the intensity-dependent field broadening. Additional studies of the transition from inhomogeneous to homogeneous broadening may initiate a further development of the theory of interatomic interactions in a high-density atomic gas.
In nanocells the dressed atomic states can be investigated in absorption spectra of the dense atomic gases SarkisyanPRL2012, SarkisyanPRL2018. The unique nonlinear properties of ultrathin layers of ”dressed dense atomic gases” might be useful for development of new quantum devices. The ”dressed states” spectroscopy may be applied to the study of many-body interactions in ultracold nonideal plasmas with the resonance transitions in the visible range ZelenerJETPL2021, ZelenerPRL2024.
@@
Vladimir Sautenkov: Conceptualization, Methodology, Validation, Writing – original draft. Sergey Saakyan: Investigation, Software, Validation, Formal analysis, Writing – review & editing. Andrei Bobrov: Formal analysis, Methodology, Writing – review & editing. Boris B. Zelener: Supervision, Resources, Writing – review & editing, Funding acquisition.